in a BaI2 solution if the concentration of I- = 0.386 M then what is the concentration of BaI2


1
Expert's answer
2021-12-15T10:17:31-0500

BaI2(aq) + Ba2+(aq) + 2 I-(aq)

According to dissociation equation, BaI2 concentration = 1/2 x I-(aq) = 1/2 x 0.386 = 0.193 M.